# Safe Output Regulation of Coupled Hyperbolic PDE-ODE Systems

### v1

@@ 4.1 State observer and disturbance estimator
where $\Lambda_{1}(x)$, $\Lambda(x)$ satisfy $q_{2}\Lambda_{1}^{\prime}(x)-\Lambda_{1}(x)S_{d}+\bar{K}_{1}(x)=0$, $\Lambda_{1}(1)=\bar{G}_{5}$, $q_{1}\Lambda^{\prime}(x)+\Lambda(x)S_{d}-\bar{K}_{2}(x)=0$, $\Lambda(0)=p\Lambda_{1}(0)+\bar{G}_{4}$, and defining $p_{1}(x)$, $p_{2}(x)$ in the observer gains (104), (105) are $p_{1}(x)=-\Lambda(x)L_{d}$, $p_{2}(x)=-\Lambda_{1}(x)L_{d}$, then (106)–(110) is converted into
@@
$\bar{\alpha}(0,t)$ $=p\bar{\beta}(0,t),$ (117)
where $L_{d}$ is chosen such that $S_{d}-L_{d}\Lambda(1)$ Hurwitz. The final target observer error system we arrive at is (113)–(117) with (96). Before presenting the results, we propose the following lemma and assumptions that will be used in establishing the result about the observer error system.
Lemma 2.
For a Hurwitz matrix $A_{\rm H}$, we have the following estimate
$\|e^{A_{\rm H}t}\|_{2}\leq M_{\rm H}(t)e^{-\sigma_{1}t},\forall t\geq 0$ (118)
where $\sigma_{1}=-\max\Re\lambda(A_{\rm H})$, and where $M_{\rm H}=\|P_{1}\|_{2}\|P_{1}^{-1}\|_{2}>0$ which is constant when $A_{\rm H}$ is diagonalizable, otherwise, $M_{\rm H}(t)=\|P_{1}\|_{2}\|P_{1}^{-1}\|_{2}\sum_{j=0}^{n_{k}}\frac{t^{j}}{j!}>0$ which is a polynomial in $t$. The matrix $P_{1}$ is invertible, satisfying
$A_{\rm H}=P_{1}\Lambda P_{1}^{-1}$ (119)
where $\Lambda$ is either a diagonal matrix (if $A_{\rm H}$ is diagonalizable) or a Jordan matrix (in which the size of the largest Jordan block is $n_{k}$).
Proof.
The proof is shown in the Appendix 0.4.
For the observer-based output-feedback control in this section, we require the following assumptions. Assumptions 3, 4 indicate the bounds of the initial values of the unmeasured PDE states and the disturbances are known, but arbitrary, which will be used in estimating the upper bounds of the observer errors required in building the output-feedback safe controller in the next subsection.
@@
The bounds of initial states are known and arbitrary, i.e., $(z[t_{0}],w[t_{0}])\in\mathcal{D}$, where $\mathcal{D}=\{(\ell_{1}(x),\ell_{2}(x))|\underline{z}_{0}(x)\leq\ell_{1}(x)\leq\overline{z}_{0}(x),\underline{w}_{0}(x)\leq\ell_{2}(x)\leq\overline{w}_{0}(x)\}$. The functions $\underline{z}_{0}(x)$, $\underline{w}_{0}(x)$, $\overline{z}_{0}(x)$, $\overline{w}_{0}(x)$, are known and arbitrary.
@@
The initial values of disturbances satisfy $v(0)\in\mathcal{D}_{v}$, where $\mathcal{D}_{v}=\{(\ell_{1},\cdots,\ell_{n_{v}})\in\mathbb{R}^{n_{v}}|\underline{v}_{i}\leq\ell_{i}(0)\leq\overline{v}_{i},i=1,2,\cdots,n_{v}\}$, especially, $v_{d}(0)\in\mathcal{D}_{d}$ where $\mathcal{D}_{d}=\{(\ell_{1},\cdots,\ell_{n_{d}})\in\mathbb{R}^{n_{d}}|\underline{v}_{i}\leq\ell_{i}(0)\leq\overline{v}_{i},i=1,2,\cdots,n_{d}\}$.
@@
For Assumption 5, the exact value of $h(e(\bar{t}_{0}),\bar{t}_{0})$ is not required but its sign that will be used in determining the form of $\sigma$ in (49), which can be judged based on the known bounds of the initial values in Assumptions 3, 4.
@@
The time-varying function $h$ is $n$ times differentiable with respect to each of its arguments, i.e., $e$ as well as $t$, and satisfies that $\frac{\partial h(e,t)}{\partial e}$ is continuous as well as $\frac{\partial h(e,t)}{\partial e}\neq 0,~\forall e\in\mathbb{R}$, $t\in[t_{0},\infty)$. Besides, $|h(e(t),t)|<\infty\Rightarrow|e(t)|<\infty$ and $\lim_{t\to\infty}h(e(t),t)=0\Rightarrow\lim_{t\to\infty}e(t)=0.$ Also, the barrier function $h$ ensures that there exists a positive $\xi_{e}$ such that
@@
Due to the uncertainties of the plant state and the external disturbances, this assumption about the barrier function $h$ is less general than Assumption 2. This assumption can be verified easily when planning the barrier function $h$ using (53) (47), (48).
@@
[…]
@@
For initial data $\hat{w}[0]\in L^{2}(0,1)$, $\hat{z}[0]\in L^{2}(0,1)$, $\hat{Y}(0)\in\mathbb{R}^{n}$, $\hat{v}_{d}(0)\in\mathbb{R}^{n_{d}}$, $\hat{v}_{r}(0)\in\mathbb{R}^{n_{r}}$, with the observer (89)–(94) where the observer gains $L_{1}(x),L_{2}(x)$ are defined by (104), (105), and $L_{d}$, $L_{r}$ satisfies that $S_{d}-L_{d}\Lambda(1)$ as well as $S_{r}-L_{r}\bar{P}_{r}$ are Hurwitz, respectively, the observer error system (96)–(101) is well-posed in the sense of $\tilde{w}[t]\in L^{2}(0,1)$, $\tilde{z}[t]\in L^{2}(0,1)$, $\tilde{Y}(t)\in\mathbb{R}^{n}$, $\tilde{v}(t)\in\mathbb{R}^{n_{v}}$ for $t\in[t_{0},\infty)$. Moreover, the exponential convergence to zero of the observer errors is achieved in the sense that
$\tilde{\Omega}_{e}\leq\Upsilon_{1}\bar{M}_{d}+(1+\Upsilon_{1})M_{d}+(1+\Upsilon_{1})M_{0}M_{r}(t)e^{-\sigma_{r}t}$
$\quad:=\rho_{e1}(t),~~~t\in[0,\frac{1}{q_{1}}+\frac{1}{q_{2}}),$ (122)
$\tilde{\Omega}_{e}\leq(1+\Upsilon_{1})(\check{M}_{d}M_{2}(t)e^{-\sigma_{e}t}+M_{0}M_{r}(t)e^{-\sigma_{r}t})$
$\quad:=\rho_{e2}(t),~~~t\in[\frac{1}{q_{1}}+\frac{1}{q_{2}},\infty)$ (123)
where the positive constants $\bar{M}_{d}$, $M_{d}$, $M_{0}$, $\check{M}_{d}$, $\Upsilon_{1}$ are defined by (E.27), (E.23), (E.21), (E.25), and (E.26), respectively, and where $M_{r}(t)$, $\sigma_{r}$ as well as $M_{e}(t)$, $\sigma_{e}$ are given by Lemma 2 with applying $A_{\rm H}=S_{d}-L_{d}\Lambda(1)$ and $A_{\rm H}=S_{r}-L_{r}\bar{P}_{r}$, respectively.
@@
The proof is given in the Appendix 0.5.

### v2

@@ 4.1 State observer and disturbance estimator
$\tilde{D}(t)=$ $\tilde{X}(t)-\bar{\Lambda}\tilde{v}_{d}(t)$ (114)
where $\Lambda_{1}(x)$, $\Lambda(x)$, $\bar{\Lambda}$ satisfy
$q_{2}\Lambda_{1}^{\prime}(x)-\Lambda_{1}(x)S_{d}+\bar{K}_{1}(x)=0,$ (115)
$\Lambda_{1}(1)=\bar{G}_{5},$ (116)
$q_{1}\Lambda^{\prime}(x)+\Lambda(x)S_{d}-\bar{K}_{2}(x)=0,$ (117)
$\Lambda(0)=p\Lambda_{1}(0)+\bar{G}_{4}+C\bar{\Lambda},$ (118)
$(A-L_{y}C_{1})\bar{\Lambda}+\bar{K}_{0}-\bar{\Lambda}S_{d}=0$ (119)
The solution of (115)–(119) is easy to obtain through the following process. Denote by $\psi_{i}$ the linearly independent eigenvectors of $S_{d}$ associated with the eigenvalues $v_{i}$, $i=1,2,\ldots,n_{d}$. Post-multiplying (115)–(119) by $\psi_{i}$ and introducing
$\iota_{i}=\Lambda\psi_{i},~\iota_{1,i}=\Lambda_{1}\psi_{i},~\bar{\iota}_{i}=\bar{\Lambda}\psi_{i},$ (120)
we obtain $q_{1}\iota^{\prime}_{i}(x)+v_{i}\iota_{i}(x)-\bar{K}_{2}(x)\psi_{i}=0$, $\iota_{i}(0)=p\Lambda_{1}(0)\psi_{i}+\bar{G}_{4}\psi_{i}+C\bar{\Lambda}\psi_{i}$, $q_{2}\iota^{\prime}_{1,i}(x)-v_{i}\iota_{1,i}(x)+\bar{K}_{1}(x)\psi_{i}=0$, $\iota_{1,i}(1)=\bar{G}_{5}\psi_{i}$, $\bar{\iota}_{i}=-((A-L_{y}C_{1})-v_{i}I)^{-1}\bar{K}_{0}\psi_{i}$.
Remark 4.
Since $A-L_{y}C_{1}$ is Hurwitz, all its eigenvalues lie strictly in the open left half-plane. On the other hand, the eigenvalues of $S_{d}$ lie on the imaginary axis, as described in Sec. 2.2. It follows that $\sigma(A-L_{y}C_{1})\cap\sigma(S_{d})=\varnothing.$ Consequently, the matrix $(A-L_{y}C_{1})-v_{i}I$ is invertible.
Therefore, according to (120), the solution of (115)–(119) are $\Lambda=[\iota_{1},\cdots,\iota_{n_{d}}]\bar{V}^{-1}$, $\Lambda_{1}=[\iota_{1,1},\cdots,\iota_{1,n_{d}}]\bar{V}^{-1}$, $\bar{\iota}=[\bar{\iota}_{1},\cdots,\bar{\iota}_{n_{d}}]\bar{V}^{-1}$, where $\bar{V}=[\psi_{1},\cdots,\psi_{n_{d}}]$, and where $\iota_{i}$ and $\iota_{1,i}$ are obtained as
$\iota_{i}(x)=e^{-\frac{v_{i}}{q_{1}}x}\bigg(pe^{\frac{-v_{i}}{q_{2}}}\bar{G}_{5}\psi_{i}+\frac{p}{q_{2}}\int_{0}^{1}e^{\frac{-v_{i}}{q_{2}}\xi}\bar{K}_{1}(\xi)\psi_{i}d\xi$
$+\bar{G}_{4}\psi_{i}-C\big((A-L_{r}C_{1})-v_{i}I\big)^{-1}\bar{K}_{0}\psi_{i}\bigg)$
$+\frac{1}{q_{1}}\int_{0}^{x}e^{-\frac{v_{i}}{q_{1}}(x-\xi)}\bar{K}_{2}(\xi)\psi_{i}d\xi$ (121)
and $\iota_{1,i}(x)=e^{\frac{-v_{i}}{q_{2}}(1-x)}\bar{G}_{5}\psi_{i}-\frac{1}{q_{2}}\int_{x}^{1}e^{\frac{-v_{i}}{q_{2}}(\xi-x)}\bar{K}_{1}(\xi)\psi_{i}d\xi.$
Lemma 2.
The pair $(\Lambda(1),S_{d})$ is observable if and only if $\tilde{N}(v_{i})\bar{P}_{d}\psi_{i}\neq 0$, $i=1,\cdots,n_{d}$, where $\tilde{N}(s)$ is the numerator of the transfer matrix from $d(t)$ to the measurement $z(1,t)$.
Proof.
Applying the same transformation of (101)–(103) (replacing $\tilde{z},\tilde{w},\tilde{Y},\tilde{\alpha},\tilde{\beta},\tilde{X}$ with $z,w,Y,\check{\alpha},\check{\beta},\check{Y}$) into the plant (1)–(5), we obtain $\dot{\check{X}}(t)=(A-L_{r}C_{1})\check{X}(t)-q_{1}K_{0}(1)\check{\alpha}(1,t)+\bar{K}_{0}v_{d}(t),$ $\check{\alpha}_{t}(x,t)=-q_{1}\check{\alpha}_{x}(x,t)-q_{1}K^{11}(x,1)\check{\alpha}(1,t)+\bar{K}_{2}(x)v_{d}(t),$ $\check{\beta}_{t}(x,t)=q_{2}\check{\beta}_{x}(x,t)-q_{1}K^{21}(x,1)\check{\alpha}(1,t)+\bar{K}_{1}(x)v_{d}(t),$ $\check{\beta}(1,t)=q\check{\alpha}(1,t)+\bar{G}_{5}v_{d}(t),$ $\check{\alpha}(0,t)=p\check{\beta}(0,t)+C\check{X}(t)+\bar{G}_{4}v_{d}(t).$ Rewriting it in the frequency domain, considering $z(1,t)=\check{\alpha}(1,t)$, the transfer function from $v_{d}(t)$ to $z(1,t)$ is obtained as $\frac{z(1,s)}{v_{d}(s)}=\frac{N(s)}{D(s)}$, where $D(s)=1+\int_{0}^{1}e^{-(s/q_{1})(1-\xi)}K^{11}(\xi,1)d\xi-pqe^{-s\left(\frac{1}{q_{1}}+\frac{1}{q_{2}}\right)}+e^{-s/q_{1}}p\int_{0}^{1}\frac{q_{1}}{q_{2}}e^{-(s/q_{2})\xi}K^{21}(\xi,1)d\xi+e^{-s/q_{1}}q_{1}C\big(sI-(A-L_{r}C_{1})\big)^{-1}K_{0}(1),$ and where
$N(s)=e^{-s/q_{1}}C\big(sI-(A-L_{r}C_{1})\big)^{-1}\bar{K}_{0}$
$+e^{-s/q_{1}}\Big(p\big(\bar{G}_{5}e^{-s/q_{2}}+\Psi_{1}(s)\big)+\bar{G}_{4}\Big)+\Phi_{2}(s),$ (122)
with $\Phi_{2}(s)=\frac{1}{q_{1}}\int_{0}^{1}e^{-\frac{s}{q_{1}}(1-\xi)}\bar{K}_{2}(\xi)d\xi$ and $\Psi_{1}(s)=\frac{1}{q_{2}}\int_{0}^{1}e^{-\frac{s}{q_{2}}\xi}\bar{K}_{1}(\xi)d\xi.$ Replacing all $\bar{G}_{i}$ involved in $N(s)$ (4.1) with $G_{i}\bar{P}_{d}$, we have $N(s)=\tilde{N}(s)\bar{P}_{d}$, where $\tilde{N}(s)=\frac{z(1,s)}{d(s)}$ is the transfer function from $d(t)$ to $z(1,t)$ considering $d(s)=\bar{P}_{d}v_{d}(s)$. By the PBH (Popov–Belevitch–Hautus) observability test, the pair $(\Lambda(1),S_{d})$ is observable if and only if $\Lambda(1)\psi_{i}\neq 0$, i.e., $\iota_{i}(1)\neq 0$ according to (120). That is, $\iota_{i}(1)=N(v_{i})\psi_{i}=\tilde{N}(v_{i})\bar{P}_{d}\psi_{i}\neq 0$ recalling (4.1).
Remark 5.
Lemma 2 shows the observability condition for the disturbance observer, i.e., all eigenmodes of the disturbance model $\dot{v}_{d}(t)=S_{d}v_{d}(t)$ can be transferred from $d(t)$ to the measurement $z(1,t)$ in steady state, which allows the disturbance observer can reconstruct the state $v_{d}(t)$ from this measurement. Note that $\bar{P}_{d}\psi_{i}\neq 0$ is ensured by Assumption 1 according to PBH observability test.
Defining $p_{1}(x)$, $p_{2}(x)$ in the observer gains (104), (105) are $p_{1}(x)=-\Lambda(x)L_{d}$, $p_{2}(x)=-\Lambda_{1}(x)L_{d}$, then (106)–(111) is converted into
@@
$\bar{\alpha}(0,t)$ $=p\bar{\beta}(0,t)+C\tilde{D}(t),$ (128)
where $L_{0}$ is chosen as
$L_{0}=\bar{\Lambda}L_{d}+\bar{P}_{1}.$ (129)
The final target observer error system we arrive at is (123)–(128) with (95). The observer gain $L_{d}$ is chosen such that $S_{d}-L_{d}\Lambda(1)$ Hurwitz and $L_{r}$ is selected such that $S_{r}-L_{r}\bar{P}_{r}$ Hurwitz, considering $(S_{d},\Lambda(1))$ and $(S_{r},\bar{P}_{r})$ are observable according to Assumption 1 and Lemma 2.
Before presenting the results, we propose the following lemma and assumptions that will be used in establishing the result about the observer error system. For the observer-based output-feedback control in this section, we require the following assumptions. Assumptions 4, 5 indicate the bounds of the initial values of the unmeasured PDE states and the disturbances are known, but arbitrary, which will be used in estimating the upper bounds of the observer errors required in building the output-feedback safe controller in the next subsection.
@@
The bounds of initial states are known, i.e., $\|z(x,0)\|^{2}\leq\overline{z}_{0}$, $\|w(x,0)\|^{2}\leq\overline{w}_{0}$, $|Y(0)|^{2}\leq\overline{Y}_{0}$.
@@
The initial values of external signals satisfy $|v(0)|^{2}\in\overline{v}_{0}$, where $\overline{v}_{0}$ is known.
@@
For Assumption 6, the exact value of $h(e(\bar{t}_{0}),\bar{t}_{0})$ is not required but its sign that will be used in determining the form of $\sigma$ in (49), which can be judged based on the known bounds of the initial values in Assumptions 4, 5.
@@
The time-varying function $h$ is $n$ times differentiable with respect to each of its arguments, i.e., $e$ as well as $t$, and satisfies that $\frac{\partial h(e,t)}{\partial e}$ is continuous as well as $\frac{\partial h(e,t)}{\partial e}\neq 0,~\forall e\in\mathbb{R}$, $t\in[t_{0},\infty)$. Besides, there exists a positive constant $\xi$ such that $|e(t)|<\xi|h(e(t),t)|$. Also, the barrier function $h$ ensures that there exists a positive $\xi_{e}$ such that
@@
Assumptions 3 and 7 are independent assumptions for the barrier function $h$ in the state-feedback and output-feedback cases, respectively. Due to the uncertainties of the plant state and the external disturbances, Assumption 7 is less general than Assumption 3. Assumption 7 can also be verified easily when planning the barrier function $h$ using (54), (47), (48).
@@
$\tilde{\Omega}_{e}=(\|\tilde{z}(\cdot,t)\|^{2}+\|\tilde{w}(\cdot,t)\|^{2}+|\tilde{v}|^{2}+|\tilde{Y}|^{2})^{\frac{1}{2}},$ (131)
@@
For initial data $\hat{w}[0]\in L^{\infty}(0,1)$, $\hat{z}[0]\in L^{\infty}(0,1)$, $\hat{Y}(0)\in\mathbb{R}^{n}$, $\hat{v}_{d}(0)\in\mathbb{R}^{n_{d}}$, $\hat{v}_{r}(0)\in\mathbb{R}^{n_{r}}$, and observability condition in Lemma 2 holds, with the observer (87)–(93) where the observer gains $L_{1}(x)$, $L_{2}(x)$, $L_{0}$ are defined by (104), (105), (129), and $L_{d}$, $L_{r}$, $L_{y}$ satisfies that $S_{d}-L_{d}\Lambda(1)$, $S_{r}-L_{r}\bar{P}_{r}$, and $A-L_{y}C_{1}$ are Hurwitz, respectively, the observer errors in the observer error system (95)–(100) are exponentially convergent to zero in the sense that
$\tilde{\Omega}_{e}\leq\Upsilon_{2}e^{-\sigma_{r}t}:=\rho_{1e}(t),~~~t\in[0,\infty)$ (132)
where $\sigma_{r}>1$ depends on the Hurwitz matrices $S_{d}-L_{d}\Lambda(1)$, $S_{r}-L_{r}\bar{P}_{r}$, $A-L_{y}C_{1}$, and where $\Upsilon_{2}$ depends on the bounds of the kernels in the transformations (101)–(103), (112)–(114), and bounds of the initial state in Assumptions 4 and 5.
@@
The estimate (132) can be obtained by the following Lyapunov analysis. Define
$V(t)=$ $\tilde{v}^{T}P_{1}\tilde{v}+{r_{a}}(\tilde{D}^{T}P_{2}\tilde{D}+\frac{b}{2}\int_{0}^{1}\bar{\beta}(x,t)^{2}e^{x}dx)$
$+\frac{r_{c}}{2}\int_{0}^{1}\bar{\alpha}(x,t)^{2}e^{-x}dx$ (133)
where the positive constants $r_{a},r_{b},r_{c}$ are to be determined later. The positive definite matrix $P_{1}={P_{1}}^{T}$ is the solution to the Lyapunov equation $A_{\rm v}^{T}P_{1}+P_{1}A_{\rm v}=-Q_{1}$ for some $Q_{1}={Q_{1}}^{T}>0$, where $A_{\rm v}=[S_{r}-L_{r}\bar{P}_{r},0;0,S_{d}-L_{d}\Lambda(1)]$. The positive definite matrix $P_{2}={P_{2}}^{T}$ is the solution to the Lyapunov equation $(A-L_{y}C_{1})^{T}P_{2}+P_{2}(A-L_{y}C_{1})=-Q_{2}$ for some $Q_{2}={Q_{2}}^{T}>0$. Defining $(\tilde{\Omega}_{T}(t)=\|\bar{\alpha}(\cdot,t)\|^{2}+\|\bar{\beta}(\cdot,t)\|^{2}+|\tilde{v}|^{2}+|\tilde{D}|^{2})^{\frac{1}{2}}$, we have
$\xi_{1}\tilde{\Omega}_{T}(t)^{2}\leq V(t)\leq\xi_{2}\tilde{\Omega}_{T}(t)^{2}$ (134)
for some positive $\xi_{1}$ and $\xi_{2}$. Taking the derivative of (133) along (123)–(128) with (95), applying Young’s and cauthy-schwarz inequalities, we get
$\dot{V}(t)\leq-\frac{1}{2}\lambda_{\rm min}(Q_{1}){{v(t)}^{2}}-\bigg(\frac{{q_{1}}{r_{c}}}{2}-\frac{{2|P_{1}B_{d}|^{2}}}{\lambda_{\rm min}(Q_{1})}\bigg)\bar{\alpha}(1,t)^{2}$
$-\frac{r_{c}}{2}q_{1}\int_{0}^{1}{{}{e^{-x}}{\bar{\alpha}}{{(x,t)}}^{2}}dx-\frac{{r_{a}}q_{2}}{2}\int_{0}^{1}{e^{x}}{\bar{\beta}}{{(x,t)}}^{2}dx$
$-(r_{a}-2r_{c}q_{1}|C|^{2})\lambda_{\rm min}(Q_{2})|\tilde{D}|^{2}$
$-(q_{2}r_{a}-2p^{2}r_{c}q_{1})\bar{\beta}(0,t)^{2}.$ (135)
Choosing $r_{c}>\frac{{4|P_{1}B_{d}|^{2}}}{q_{1}\lambda_{\rm min}(Q_{1})}$, $r_{a}>2r_{c}\max\{2q_{1}|C|^{2},p^{2}\frac{q_{1}}{q_{2}}\}$, where $B_{d}=[0,L_{d}]^{T}$, we arive at $\dot{V}=-\sigma_{v}V(t)$, where $\sigma_{v}=\frac{1}{2\xi_{2}}\min\{\lambda_{\rm min}(Q_{1}),{r_{c}}q_{1}e^{-1},{{r_{a}}q_{2}},{r_{a}}\lambda_{\rm min}(Q_{2})\}$. Therefore, $\tilde{\Omega}_{T}(t)\leq\sqrt{\frac{\xi_{2}}{\xi_{1}}}\tilde{\Omega}_{T}(0)e^{-\sigma_{r}t}$, where $\sigma_{r}=\frac{1}{2}\sigma_{v}$. Estimate (132) is thus obtained by applying the transformations (101)–(103), (112)–(114), and bounding the initial values by Assumptions 4 and 5, i.e., $\tilde{\Omega}_{e}(0)\leq(\overline{z}_{0}+\overline{w}_{0}+\overline{v}_{0}+\overline{Y}_{0})^{\frac{1}{2}}$.